# Boosting Generalization in Diffusion-Based Neural Combinatorial Solver via Inference Time Adaptation

### v2

@@ 1 Introduction
Figure 1: Overview of Inference Time Adaptation for achieving zero-shot cross-problem generalization, shifting from pre-trained problem $G$ (TSP) solution space to target problem $G^{\prime}$ (PCTSP).
Figure 2: The proposed Inference Time Adaptation framework. This approach combines (1) energy-guided sampling, which incorporates problem-specific objectives and constraints, with (2) a recursive denoising-decoding travel strategy for solution refinement, enabling zero-shot cross-problem transfer without training. The Optimality Gap ($\downarrow$) on PCTSP-20 is reduced from $19.21\%$ to $4.20\%$.
This work introduces an inference time adaptation framework (illustrated in Figure 2) designed to enhance the generalization capabilities of diffusion-based Neural Combinatorial Optimization (NCO) solvers without additional training costs. We scale up the inference time by combining two key components: (1) energy-guided sampling, which integrates problem-specific objective functions and constraints into the denoising process, and (2) a recursive denoising-decoding strategy applied at intermediate noise levels. This inference time adaptation framework enables zero-shot cross-problem transfer while maintaining solution feasibility. Furthermore, we present a theoretical analysis to deepen the understanding of the problem transfer capability afforded by this inference time adaptation when applied to pre-trained models. Through experimental evaluation of the TSP-trained diffusion model on more complex variants, the Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP), we empirically demonstrate its effective zero-shot transferability across different problems of increasing complexity, while maintaining consistent performance across different problem scales. Our work represents a significant step toward more flexible and generalizable diffusion-based combinatorial optimization solvers, reducing the need for problem-specific model training while maintaining competitive performance.
@@ 2 Related Works
Neural Network-based Combinatorial Solvers. Neural Combinatorial Optimization (NCO) approaches focus on leveraging neural networks to learn feasible solution distributions for combinatorial optimization problems [3, 18]. Autoregressive construction solvers [9, 10, 19, 20, 21] are built upon the success of transformer-based [22] and reinforcement learning architectures in sequential generation tasks. However, non-autoregressive construction solvers [23, 24, 25, 11, 6, 26] have also been proposed to learn high-quality solution distributions.
Diffusion-based Generative Modeling. Recent advances in generative modeling have revolutionized various domains through diverse approaches, including Variational Autoencoders (VAE) [27], Generative Adversarial Networks (GAN) [28], Diffusion models [4], and GFlowNet [29]. In particular, score-based diffusion models [4, 5, 30, 31, 32, 33] have emerged as a powerful framework operating in continuous domains.
Beyond their state-of-the-art performance in traditional generative tasks, these models have shown remarkable potential in combinatorial optimization (CO). Pioneering work by [6] established new state-of-the-art results for the Traveling Salesman Problem (TSP) by adapting discrete diffusion models [34] to graph structures. Building upon this foundation, [7] enhanced the framework’s performance through gradient search and the current state-of-the-art [8] introduced consistency models to achieve single-step generation. [26] proposed the first diffusion-based unsupervised learning framework.
Training-free Guidance for Diffusion Models. Conditional generation has emerged as a crucial component in real-world applications, enabling precise control over generated outputs. While traditional approaches such as classifier guidance [32] and classifier-free guidance [35] have proven effective, they require substantial computational overhead due to additional training requirements for either the classifier or the diffusion model.
A promising alternative has emerged through training-free guidance methods [13, 14, 15, 16], which are guided by pre-trained networks or loss functions. In the context of discrete diffusion models, this approach remained largely unexplored until [7] pioneered the adaptation of loss-based guidance during inference, building upon the pre-trained discrete diffusion solver framework [6]. Our work demonstrates the significant potential of modifying the infernce phase in enhancing the zero-shot cross-problem transfer of diffusion-based NCO solvers.
@@ 3.1 Graph-based CO Problems
Combinatorial optimization (CO) problems on graphs are fundamental to numerous real-world applications. Following recent advances [6, 7], we address these problems by formalizing graph-based CO instances. Following [6, 7, 8], we represent each problem instance as an undirected graph $G(V,E)\in\mathcal{G}$, where $V$ and $E$ denote the vertex and edge sets, respectively. This representation encompasses both vertex selection and edge selection problems, covering a broad spectrum of practical CO scenarios. For any instance $G\in\mathcal{G}$, we define a binary decision variable $\mathbf{x}\in\mathcal{X}_{\mathcal{G}}$, where $\mathcal{X}_{\mathcal{G}}=\{0,1\}^{N}$ represents the feasible solution space. The optimization objective is to find the optimal solution $\mathbf{x}^{*}$ that minimizes a problem-specific objective function $\phi(\cdot;G):\{0,1\}^{N}\rightarrow\mathbb{R}$:
@@
Here $f_{\text{cost}}(\cdot;G)$ measures the solution quality, and $g_{i}(\cdot;G)$ enforces problem-specific constraints through a penalty coefficient $\mu>0$. The validity functions $g_{i}(\cdot;G)$ returns 00 for feasible solutions and is strictly positive for infeasible ones.
3.2 Probabilistic Modeling for CO Problems
@@
$\mathcal{E}(\cdot;G):=\rho(y,\phi(\cdot;G)),$ (3)
@@
$p(y|\mathbf{x};G)=\frac{\exp\big{(}-\frac{1}{\tau}\mathcal{E}(y,\mathbf{x};G)\big{)}}{\mathcal{Z}},\quad\text{where}\;\mathcal{Z}=\sum_{\mathbf{x}}\exp\Big{(}-\frac{1}{\tau}\mathcal{E}(y,\mathbf{x};G)\Big{)},$ (4)
@@
3.3 Discrete Diffusion Generative Models
We adopt a discrete diffusion framework [34] to effectively sample optimal solutions from the learned distribution $p_{\theta}(\mathbf{x}|y^{*},G)$. In contrast to continuous diffusion models that employ Gaussian noise, the discrete formulation is particularly well-suited for graph-based combinatorial optimization problems [6, 7, 8].
The diffusion process consists of two key components: a forward process that gradually corrupts the data, and a reverse process that learns to reconstruct the original distribution. The forward process $q(\mathbf{x}_{1:T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ maps clean data $\mathbf{x}_{0}\sim q(\mathbf{x}_{0}|G)$ to a sequence of increasingly corrupted latent variables $\mathbf{x}_{1:T}$. The reverse process $p_{\theta}(\mathbf{x}_{0:T}|G)=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},G)$ learns to gradually denoise these latent variables to recover the original distribution. From a variational perspective, we optimize the model by minimizing an upper bound on the negative log-likelihood, where $C$ is a constant:
$L(\theta)$ $=\mathbb{E}_{G\sim\mathcal{G}}[-\log p_{\theta}(\mathbf{x}_{0}|G)]$ (6)
$\leq\sum_{t=2}^{T}\mathbb{E}_{q(\mathbf{x}_{t}|\mathbf{x}_{0})}\bigg{[}D_{KL}[q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})||p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},G)]-\log p_{\theta}(\mathbf{x}_{0}|\mathbf{x}_{1},G)\bigg{]}+C.$
For discrete state spaces, we define the forward process using a categorical distribution:
@@ 3.3 Discrete Diffusion Generative Models
where $[\mathbf{Q}_{t}]_{ij}$ denotes the state transition probability from state $i$ to state $j$. The $t$-step marginal distribution and posterior can be derived as:
@@
$q(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{x}_{0})$ $=\text{Cat}\left(\mathbf{x}_{t-1};\mathbf{p}=\frac{\mathbf{\widetilde{x}}_{t}\mathbf{Q}_{t}^{\top}\odot\mathbf{\widetilde{x}}_{0}\overline{\mathbf{Q}}_{t-1}}{\mathbf{\widetilde{x}}_{0}\overline{\mathbf{Q}}_{t}\mathbf{\widetilde{x}}_{t}^{\top}}\right),$
@@
A decoding strategy is essential to construct feasible solutions from the generated adjacency matrix, which is often represented as a heatmap from which an optimal solution is decoded. Among such strategies, greedy decoding is a straightforward approach that iteratively selects nodes or edges with the highest probabilities from this heatmap until a termination condition is met (e.g., forming a complete cycle in Traveling Salesperson Problem (TSP) applications).
4 Proposed Methods: Inference Time Adaptation
@@ 4.1 Energy-guided Sampling for Problem Transfer
Figure 3: Visualization of the recursive denoising-decoding travel strategy. This approach applies re-noising and energy-guided denoising within an efficient intermediate stage (Time Stage 2), based on the observation that the efficacy of energy guidance varies across different noise levels.
Let $\mathcal{G}^{\prime}=\{G^{\prime}_{i}\}_{i=1}^{n}$ denote a set of new problem instances which differs from the pre-trained problem set $\mathcal{G}$. For a new instance $G^{\prime}$ with its optimal solution pair $(\mathbf{x},y_{G^{\prime}}^{*})$, we need to estimate the new reverse process $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y_{G^{\prime}}^{*},G^{\prime})$ according to (10). Following the score estimation perspective of diffusion processes [5], we decompose the conditional score function at time step $t$ into two components:
$\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}|y^{*}_{G^{\prime}},G^{\prime})}_{\text{Posterior Score}}$ $=\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t}|G^{\prime})}_{\text{Pre-trained Prior Score}}+\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})}_{\text{Energy-guided Score}}.$ (11)
From the Bayesian perspective, $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ can be understood as the prior, which contains knowledge of the pre-trained problems (i.e. TSP problem in our experimental settings), and $p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ corresponds to the adjustive likelihood that incorporates additional constraints or objectives of the variant problem. We sample from the posterior $p_{\theta}(\mathbf{x}_{t}|y^{*}_{G^{\prime}},G^{\prime})$ to generate high-quality solutions to the new problems.
@@
$\propto-\nabla_{\mathbf{x}_{t}}\rho(y^{*}_{G^{\prime}},\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})),$
where $\rho(y^{*}_{G^{\prime}},\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}))$ measures the distance of energy states between the optimal value and the predicted solution. Here, $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t})$ represents the predicted clean sample from the current noisy state $\mathbf{x}_{t}$, which is estimated by the logits of $N$ independent Bernoulli samples $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t})=\mathbb{E}_{\widetilde{\mathbf{x}}_{0}\sim p_{\theta}(\widetilde{\mathbf{x}}_{0}|\mathbf{x}_{t})}[\widetilde{\mathbf{x}}_{0}]$. In this work, we define $\rho(y^{*}_{G^{\prime}},\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})):=\|y^{*}_{G^{\prime}}-\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})\|_{2}$. From Eq (12), we have:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ $\propto-\nabla_{\mathbf{x}_{t}}\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}),$ (13)
$p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ $=\exp\left(-\frac{1}{\tau}\nabla_{\mathbf{x}_{t}}\phi\left(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}\right)\right).$
where $\tau$ is the energy guidance temperature to control the strength of transferring. Combining Eq. (11) and Eq. (13), we derive an energy-guided reverse sampling process for problem transfer:
$\underbrace{p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y^{*}_{G^{\prime}},G^{\prime})}_{\text{Transfer Posterior}}\propto\underbrace{p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},G^{\prime})}_{\text{Pre-trained Prior}}\underbrace{p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})}_{\text{Guidance Likelihood}}.$ (14)
In this work, we have the following log-barrier formulations for PCTSP and OP as the guided gradient:
Proposition 4.1 (Prize Collecting TSP (PCTSP)).
Given a complete graph $G=(V,E)$ with edge weights $w:E\rightarrow\mathbb{R}^{+}$, vertex prizes $r:V\rightarrow\mathbb{R}^{+}$, penalties $p:V\rightarrow\mathbb{R}^{+}$, and prize threshold $R$, find $\mathbf{x}\in\{0,1\}^{|E|}$, $\mathbf{y}\in\{0,1\}^{|V|}$ that minimizes:
@@
Proposition 4.2 (Orienteering Problem (OP)).
Given a complete graph $G=(V,E)$ with edge weights $w:E\rightarrow\mathbb{R}^{+}$, vertex scores $s:V\rightarrow\mathbb{R}^{+}$, and budget $B$, find $\mathbf{x}\in\{0,1\}^{|E|}$, $\mathbf{y}\in\{0,1\}^{|V|}$ that minimizes:
@@
4.2 Recursive Denoising-decoding Travel Strategy
Our initial experiments applying energy-guided denoising for cross-problem transfer highlighted a significant challenge: this approach alone often showed insufficient to effectively bridge the gap between the learned distribution of pre-trained problem types and the target distribution of new problem instances. This distributional divergence frequently hindered the generation of high-quality, feasible solutions for cross-problem transfer.
To address this, we first scaled up the inference time and explored a iterative rewrite strategy, adapting from [7]. This involved taking an initial solution $x_{0}^{(1)}$ for the pre-trained problem instance $G$, adding noise to obtain a highly noisy graph $x_{T}^{(1)}$, and then denoising to obtain the clean state $x_{0}^{(2)}$, aiming to progressively shift the solution towards the characteristics of the target problem. While this recursive strategy demonstrated improved transfer capabilities, it came at the cost of a considerable increase in inference time due to the repeated denoising iterations.
This led us to investigate the efficiency of the guidance mechanism itself. Inspired by the time-travel strategies in vision tasks [15], which revealed that guidance effects are not uniform across diffusion time steps, we hypothesized a similar phenomenon in NCO task. Our empirical analysis confirmed this: the impact of energy guidance varied significantly with the noise level. Crucially, we observed that guidance was most potent and beneficial at intermediate noise levels from 200 to 500 steps. In contrast, its effect was minimal on highly noisy graphs (analogous to an early time stage) and became largely redundant for nearly converged solutions (a late, refinement stage), where the solution structure is already well-defined.
Based on this observation that the impact of energy guidance at different noise levels, we developed our optimized recursive denoising-decoding travel strategy in Figure 3. Instead of applying full re-noising and iterative energy-guided denoising across many steps, our refined approach reduces the re-noising process to only a few steps. More importantly, the computationally intensive energy-guided denoising is applied for just a single step, precisely targeted at these effective intermediate noise levels. This targeted acceleration reduces the computational overhead associated with the recursive process, yielding a 5-10x speedup in inference time compared to the original iterations. This optimization not only maintains but enhances solution quality in several instances, establishing a significantly more favorable performance-to-speed trade-off.

### v3

# Boosting Cross-problem Generalization in Diffusion-Based Neural Combinatorial Solver via Inference Time Adaptation

@@ 1 Introduction
Figure 1: The proposed Inference Time Adaptation framework. This approach combines (1) energy-guided sampling, which incorporates problem-specific objectives and constraints, with (2) a recursive renoising-denosing travel for solution refinement, enabling zero-shot cross-problem transfer without training. The Optimality Gap ($\downarrow$) on PCTSP-20 is reduced from $19.21\%$ to $4.20\%$.
@@
This work introduces an inference time adaptation framework (illustrated in Figure 1) designed to enhance the generalization capabilities of diffusion-based Neural Combinatorial Optimization (NCO) solvers without additional training costs. We scale up the inference time by combining two key components: (1) energy-guided sampling and (2) a recursive renoising-denosing travel modelled as a Guided Langevin Dynamics. Our DIFU-Ada framework enables zero-shot cross-problem transfer while maintaining solution feasibility. Through experimental evaluation of the TSP-trained diffusion model on more complex variants, the Prize Collecting TSP (PCTSP) and the Orienteering Problem (OP), we empirically demonstrate its effective zero-shot transferability across different problems of increasing complexity, while maintaining consistent performance across different problem scales. Our work represents a significant step toward more flexible and generalizable diffusion-based combinatorial optimization solvers, reducing the need for problem-specific model training while maintaining competitive performance.
@@ 2 Related Works
Neural Network-based Combinatorial Solvers. Neural Combinatorial Optimization (NCO) leverages neural networks to learn solution distributions for complex optimization problems [3, 18]. Prominent approaches include autoregressive solvers [9, 10, 19, 20, 21] and non-autoregressive methods [22, 23, 24, 11, 6, 25].
Diffusion-based Generative Modeling. Score-based diffusion models [4, 5, 26, 27, 28, 29] have become a leading generative framework. Their application to CO was pioneered by [6] for the Traveling Salesman Problem and has been rapidly advanced by methods incorporating gradient search [7], single-step consistency models [8], and unsupervised frameworks [25].
Training-free Guidance for Diffusion Models. To control generation, training-free guidance methods [13, 14, 15, 16] offer an efficient alternative to training-intensive classifier [28] or classifier-free [30] approaches. This strategy was first adapted for discrete diffusion solvers by [7], building on the framework from [6]. Our work further demonstrates that modifying the inference phase is a powerful method for enhancing the zero-shot cross-problem transfer of diffusion-based NCO solvers.
@@ 3.1 Graph-based CO Problems
Combinatorial optimization (CO) problems on graphs are fundamental to numerous real-world applications. Following recent advances [6, 7], we address these problems by formalizing graph-based CO instances. We represent each problem instance as an undirected graph $G(V,E)\in\mathcal{G}$, where $V$ and $E$ denote the vertex and edge sets, respectively. This representation encompasses both vertex selection and edge selection problems, covering a broad spectrum of practical CO scenarios. For any instance $G\in\mathcal{G}$, we define a binary decision variable $\mathbf{x}\in\mathcal{X}_{\mathcal{G}}$, where $\mathcal{X}_{\mathcal{G}}=\{0,1\}^{N}$ represents the feasible solution space. The optimization objective is to find the optimal solution $\mathbf{x}^{*}$ that minimizes a problem-specific objective function $\phi(\cdot;G)\mathrel{\mathop{\ordinarycolon}}\{0,1\}^{N}\rightarrow\mathbb{R}$:
@@
Here $f_{\text{cost}}(\cdot;G)$ measures the solution quality, and $g_{i}(\cdot;G)$ enforces problem-specific constraints through a penalty coefficient $\mu>0$. The validity functions $g_{i}(\cdot;G)$ returns 0 for feasible solutions and is strictly positive for infeasible ones.
3.2 Energy-based Probabilistic Modeling
@@
$\mathcal{E}(\cdot;G)\mathrel{\mathop{\ordinarycolon}}=\rho(y,\phi(\cdot;G)),$ (3)
@@
$p(y|\mathbf{x};G)=\frac{\exp\big{(}-\frac{1}{\tau}\mathcal{E}(y,\mathbf{x};G)\big{)}}{\mathcal{Z}},$ (4)
where $\mathcal{Z}=\sum_{\mathbf{x}}\exp\Big{(}-\frac{1}{\tau}\mathcal{E}(y,\mathbf{x};G)\Big{)},$ (5)
@@ 3.3 Discrete Diffusion Generative Models
We adopt a discrete diffusion generative model [34] to effectively sample optimal solutions from the learned distribution $p_{\theta}(\mathbf{x}|y^{*},G)$ [6, 7, 8].
The diffusion process consists of two key components: a forward process that gradually corrupts the data, and a reverse process that learns to reconstruct the original distribution. The forward process $q(\mathbf{x}_{1\mathrel{\mathop{\ordinarycolon}}T}|\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}|\mathbf{x}_{t-1})$ maps clean data $\mathbf{x}_{0}\sim q(\mathbf{x}_{0}|G)$ to a sequence of increasingly corrupted latent variables $\mathbf{x}_{1\mathrel{\mathop{\ordinarycolon}}T}$. The reverse process $p_{\theta}(\mathbf{x}_{0\mathrel{\mathop{\ordinarycolon}}T}|G)=p(\mathbf{x}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},G)$ learns to gradually denoise these latent variables to recover the original distribution. For discrete state spaces, we define the forward process using a categorical distribution:
@@
where $\mathbf{Q}_{t}]_{ij}$ denotes the state transition probability from state $i$ to state $j$. The $t$-step marginal distribution and posterior can be derived as:
@@
4 Method: Inference Time Adaptation
@@ 4.1 Energy-guided Sampling for Problem Transfer
Let $\mathcal{G}^{\prime}=\{G^{\prime}_{i}\}_{i=1}^{n}$ denote a set of new problem instances which differs from the pre-trained problem set $\mathcal{G}$. For a new instance $G^{\prime}$ with its optimal solution pair $(\mathbf{x},y_{G^{\prime}}^{*})$, we need to estimate the new reverse process $p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},y_{G^{\prime}}^{*},G^{\prime})$ according to (10). Following the score estimation perspective of diffusion processes [5, 28], the reverse sampling for new problem types $G^{\prime}$ can be defined in a SDE formulation:
$d\mathbf{x}=\left[-\mathbf{f}(\mathbf{x},t)+g^{2}(t)\hat{s}_{\theta}(\mathbf{x},t,G^{\prime})\right]dt^{\prime}+g(t)d\mathbf{w}$ (11)
where $t=T-t^{\prime}$. We decompose the conditional score function $\hat{s}_{\theta}(\mathbf{x},t,G^{\prime})$ at time step $t$ into two components:
$\hat{s}_{\theta}(\mathbf{x},t,G^{\prime})\mathrel{\mathop{\ordinarycolon}}=\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}|y^{*}_{G^{\prime}},G^{\prime})}_{\text{Posterior Score}}=\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t}|G^{\prime})}_{\text{Pre-trained Prior Score}}+\underbrace{\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})}_{\text{Energy Potential}}.$ (12)
From the Bayesian perspective, $p_{\theta}(\mathbf{x}_{t}|G^{\prime})$ can be understood as the prior, which contains knowledge of the pre-trained problems (i.e. TSP problem in our experimental settings), and $p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ corresponds to the Energy Potential that incorporates additional constraints or objectives of the variant problem. We sample from the Posterior Score $\hat{s}_{\theta}(\mathbf{x},t,G^{\prime})$ to generate high-quality solutions to the new problem $G^{\prime}$. A theoretical analysis of how the information contained in pre-trained models can benefit the guided sampling is provided in Appendix C.
@@
where $\rho(y^{*}_{G^{\prime}},\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}))$ measures the distance of energy states between the optimal value and the predicted solution. Here, $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t})$ represents the predicted clean sample from the current noisy state $\mathbf{x}_{t}$, which is estimated by the logits of $N$ independent Bernoulli samples $\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t})=\mathbb{E}_{\widetilde{\mathbf{x}}_{0}\sim p_{\theta}(\widetilde{\mathbf{x}}_{0}|\mathbf{x}_{t})}[\widetilde{\mathbf{x}}_{0}]$. In this work, we define $\rho(y^{*}_{G^{\prime}},\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}))\mathrel{\mathop{\ordinarycolon}}=\|y^{*}_{G^{\prime}}-\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})\|_{2}$. From Eq (13), we have:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(y^{*}_{G^{\prime}}|\mathbf{x}_{t},G^{\prime})$ $\propto-\nabla_{\mathbf{x}_{t}}\rho(y^{*}_{G^{\prime}},\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime})),$ (14)
$\propto-\nabla_{\mathbf{x}_{t}}\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x}_{t});G^{\prime}),$
Now we obtain the Energy Potential by just taking the gradient of the problem-specific objective function $\phi(\widetilde{\mathbf{x}}_{0};G)$. Given a pre-trained score-based diffusion model $s_{\theta}(\mathbf{x},t,G)\approx\nabla_{\mathbf{x}_{t}}\log p_{\theta}(\mathbf{x}_{t}|G^{\prime})$,the reverse process is guided by Energy Potential modifying the score estimate as follows:
$d\mathbf{x}=\bigg{[}$ $-\mathbf{f}(\mathbf{x},t)+g(t)^{2}\bigg{(}\underbrace{\nabla_{\mathbf{x}}\log p_{\theta}(\mathbf{x}|G^{\prime})}_{\text{Pre-trained Score}}-\underbrace{\tau\nabla_{\mathbf{x}}\phi(\widetilde{\mathbf{x}}_{0}(\mathbf{x});G^{\prime})}_{\text{Energy Potential}}\bigg{)}\bigg{]}dt\textquoteright+g(t)d\mathbf{w}$ (15)
where $t=T-t^{\prime}$. For discrete diffusion, each step of the reverse process involves sampling the next state, $\mathbf{x}_{t}$, from a multivariate Bernoulli distribution. This distribution is parameterized by the output of the guided posterior $p_{\theta}(\mathbf{x}|y^{*}_{G^{\prime}},G^{\prime})$, following a standard sampling framework like DDIM [29].
Figure 2: Overview of recursive renoising-denoising travel in Inference Time Adaptation for achieving zero-shot cross-problem generalization, sequentially shifting from pre-trained problem $G$ (TSP) solution distribution to target problem distribution $G^{\prime}$ (PCTSP).
In this work, we have the well-established log-barrier formulations for PCTSP (Definition 4.1) and OP (Definition 4.2) as the plug-and-play Energy Potential following [32]:
Definition 4.1 (Energy Potential for Prize Collecting TSP (PCTSP)).
Given a complete graph $G=(V,E)$ with edge weights $w\mathrel{\mathop{\ordinarycolon}}E\rightarrow\mathbb{R}^{+}$, vertex prizes $r\mathrel{\mathop{\ordinarycolon}}V\rightarrow\mathbb{R}^{+}$, penalties $p\mathrel{\mathop{\ordinarycolon}}V\rightarrow\mathbb{R}^{+}$, and prize threshold $R$, find $\mathbf{x}\in\{0,1\}^{|E|}$, $\mathbf{y}\in\{0,1\}^{|V|}$ that minimizes:
@@
Definition 4.2 (Energy Potential for Orienteering Problem (OP)).
Given a complete graph $G=(V,E)$ with edge weights $w\mathrel{\mathop{\ordinarycolon}}E\rightarrow\mathbb{R}^{+}$, vertex scores $s\mathrel{\mathop{\ordinarycolon}}V\rightarrow\mathbb{R}^{+}$, and budget $B$, find $\mathbf{x}\in\{0,1\}^{|E|}$, $\mathbf{y}\in\{0,1\}^{|V|}$ that minimizes:
@@
4.2 Decoding from Generated Heatmaps
Distinguished from auto-regressive solvers like transformer, diffusion-based solvers for combinatorial optimization learn to generate a time-dependent matrix of probabilities, often visualized as an $N\times N$ adjacency matrix (heatmap). This heatmap represents the likelihood of each element (e.g., an edge in a graph) being part of the final solution, rather than the discrete solution itself. Consequently, a decoding strategy is essential to translate this probabilistic output into a feasible, binary solution structure.
A common and direct approach is greedy decoding, which iteratively constructs a solution by selecting the elements with the highest probabilities from the heatmap until a valid termination condition is met (e.g., forming a complete tour in the Traveling Salesperson Problem). Alternatively, the generated heatmaps can be post-processed by more sophisticated improvement heuristics, such as 2-opt [35, 7] or Monte Carlo Tree Search [6]. To ensure a fair comparison across all methods in our experiments, we apply greedy decoding.
4.3 Recursive Renoising-denoising Travel
Our preliminary experiments revealed that applying energy-guided sampling alone is often insufficient for zero-shot, cross-problem transfer. The distributional divergence between the source problem ($G$) and the target problem ($G^{\prime}$) can hinder the generation of high-quality, feasible solutions. To bridge this gap, we propose 2-phase Inference-Time Adaptation framework combining energy-guided sampling and recursive renoising-denoising travel, inspired by recent work [36, 7].
We theoretically interpret our Inference-time Adaptation as a Guided Langevin Dynamics [27, 37] process in Figure 2. It is designed to iteratively transport the solution particle from a target distribution $G$ towards a new problem distribution $G^{\prime}$. Given a solution $x^{k}$ in iteration $k$, the recursive update to next iteration $k+1$ is a discrete-time sequence of SDE:
$\mathbf{x}^{(k+1)}\leftarrow\mathbf{x}^{(k)}-\left(s_{\theta}(\mathbf{x}^{(k)},G)-\tau\nabla_{\mathbf{x}^{(k)}}\phi(\mathbf{x}^{(k)};G^{\prime})\right)+\sigma_{k}\mathbf{z},$ (16)
where $\mathbf{z}$ represent renoising. A naive implementation approach would be to recursively simulate the full SDE (i.e., repeated full re-noising and denoising) to progressively adapt the solution to the target distribution. However, this incurs a prohibitive computational cost.
Our key insight is that the energy guidance does not need to be fully applied at every step. Based on this, we implement the recursive process to only a few renoising steps and a single step denoising in each iteration, achieving a 5-10x inference speedup compared to the full recursive approach.